OpenStudy (anonymous):

11. Use w and z to solve. z = (- 5 √3 )/ 2 + (5/2)i w = 1 + (√3) i a. Convert z and w to polar form. b. Convert zw using De Moivre’s Theorem. c. Calculate z / w using De Moivre’s Theorem.

mathslover (mathslover):

First of all. We shall solve a) i) Converting z into polar form : If a + bi is the complex number then \(r\cos \theta = a\) and \(r\sin \theta = b\) Similarly, we have here : \(z = \cfrac{-5\sqrt{3}}{2} + \cfrac{5}{2} i \) as the complex number. Can you tell me what is \(r\cos \theta\) and \(r \sin \theta \) here ... ?

OpenStudy (anonymous):

juicy question. i'll leave this to u bro

OpenStudy (anonymous):

im not sure

mathslover (mathslover):

You there @kaylalynn ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i said im not sure. :(

mathslover (mathslover):

Oh;. I didn't notice that.

mathslover (mathslover):